# Testing Monotonicity in a Finite Population

### v1

@@ Abstract.
We consider the extent to which we can learn from a completely randomized experiment whether everyone has treatment effects that are weakly of the same sign, a condition we call monotonicity. From a classical sampling perspective, it is well-known that monotonicity is untestable. By contrast, we show from the design-based perspective—in which the units in the population are fixed and only treatment assignment is stochastic—that the distribution of treatment effects in the finite population (and hence whether monotonicity holds) is formally identified. We argue, however, that the usual definition of identification is unnatural in the design-based setting because it imagines knowing the distribution of outcomes over different treatment assignments for the same units. We thus evaluate the informativeness of the data by the extent to which it enables frequentist testing and Bayesian updating. We show that frequentist tests can have nontrivial power against some alternatives, but power is generically limited. Likewise, we show that there exist (non-degenerate) Bayesian priors that never update about whether monotonicity holds. We conclude that, despite the formal identification result, the ability to learn about monotonicity from data in practice is severely limited.
@@ 1. Introduction
By contrast, the design-based perspective views the $n$ units and their potential outcomes as fixed or conditioned upon, with statistical uncertainty arising only from the stochastic assignment of treatment. The data-generating process $(Y_{T},Y_{U})\sim f_{\theta}$ is then parameterized by $\theta=(\theta_{at},\theta_{nt},\theta_{d},\theta_{c})$, where $\theta_{t}$ is the count of how many of the $n$ units are of type $t$. The design-based monotonicity assumption is then that $\min\{\theta_{c},\theta_{d}\}=0$. In sum, the classical monotonicity assumption asks whether there are any units with opposite-signed treatment effects in the superpopulation from which the sample is drawn, while the design-based version asks whether there are any opposite-signed treatment effects among the $n$ units at hand.
It is well-known that data are not informative about whether monotonicity holds in the superpopulation. Suppose that we have a completely randomized experiment in which $n_{1}$ of the $n$ units are randomly assigned to treatment. The researcher then observes $(Y_{T},Y_{U})=(\sum_{i}D_{i}Y_{i},\sum_{i}(1-D_{i})Y_{i})$, summarizing how many units in each treatment group have $Y_{i}=1$. From the superpopulation perspective, we have that $(Y_{T},Y_{U})\sim g_{p}$ (where $g_{p}$ is is the push-forward of $g_{p}^{*}$ under random assignment with probability $n_{1}/n$). However, it is straightforward to show that $p$ is not point-identified (see, e.g., heckman_making_1997). In particular, for any $p_{1}$ that violates monotonicity, there exists a $p_{0}$ satisfying monotonicity such that $g_{p_{1}}=g_{p_{0}}$. Hence, from the superpopulation perspective, the data provide no information as to whether monotonicity holds. [^2] We note that monotonicity may be testable when combined with other assumptions. For example, there are tests of the joint assumptions of instrument monotonicity, exclusion, and independence (e.g. kitagawa_test_2015).
By contrast, we show that from the design-based perspective, the type counts $\theta$ are in fact identified according to the usual technical definition of identification: the mapping from the parameter to the distribution of the observed data, $\theta\mapsto f_{\theta}$, is one-to-one. At first blush, the typical definition of identification appears to suggest that the data are informative about whether the design-based version of monotonicity holds.
@@
From the frequentist perspective, we show that there exist tests of the design-based monotonicity assumption with non-trivial power against some alternatives, but we generally expect the power of these tests to be poor. In particular, we show that any test of monotonicity must have trivial power against some alternative. Moreover, we show that any test that has power against some alternative $\theta_{1}$ will have poor power at alternatives “close” to $\theta_{1}$. We formalize this by deriving upper bounds on the weighted average power (WAP) of tests in a neighborhood of $\theta_{1}$, where the weights are proportional to the frequencies expected under sampling types from a superpopulation (and thus concentrate around $\theta_{1}$, in terms of relative frequencies, when $n$ is reasonably large). We show that regardless of sample size, the WAP of a size-$\alpha$ test is no greater than $2.51\alpha$; thus, for a 5% test, the WAP is never more than 12.6%. The bound can be even tighter given a particular $n$ and a lower bound on the number of defiers/compliers. For example, if $n=100$ and compliers and defiers are each at least 5% of the finite population, then WAP is never greater than 5.06%. We also derive upper bounds on power for unbiased tests (i.e. those with power at least $\alpha$ for all alternatives). This upper bound converges to $\alpha$ as $n$ grows large, indicating that with large $n$ all unbiased tests have nearly trivial power.
@@
Taken together, our results suggest that frequentist tests provide limited learning about monotonicity. In particular, these tests can only be useful if we are interested in testing one particular alternative and not nearby ones. For example, with $n=30$ and $n_{1}=15$, the most powerful 5% test against any alternative rejects with probability 31% when there are 18 defiers and 12 compliers. However, this test never rejects when there are 17 defiers and 13 compliers. Our theoretical results show the low power of this test against neighboring alternatives is a generic feature of tests against monotonicity. We suspect that in most practical settings, researchers will not be interested in ultra-specific alternatives, and thus such frequentist tests will have relatively little use in practice.
@@
From the Bayesian perspective, we show that there do exist Bayesians who update about the probability that the design-based monotonicity assumption holds, but there also exist Bayesians who never update. The existence of a Bayesian who updates follows immediately from the fact that $\theta$ is identified. Consider a Bayesian with a two-point prior on $\theta_{0}$ satisfying monotonicity and $\theta_{1}$ violating monotonicity. Since $f_{\theta_{0}}\neq f_{\theta_{1}}$, the Bayesian updates about the probability that the null is true. However, we show that there exists a non-trivial prior $\pi^{*}$ which never updates on the probability that the null hypothesis holds, i.e. $\pi^{*}(\theta\in\Theta_{0}\mid(Y_{U},Y_{T}))=\pi^{*}(\theta\in\Theta_{0})\in(0,1)$ for all $Y_{U},Y_{T}$, where $\Theta_{0}$ denotes the set of parameter values satisfying monotonicity. [^3] Note that since the likelihood is not flat, $f_{\theta_{1}}\neq f_{\theta_{0}}$ for any $\theta_{1}\neq\theta_{0}$, this Bayesian does update their posterior about $\theta$. However, they do not update about the event $\theta\in\Theta_{0}$. Our results thus suggest that the data may be informative to some audience members, who are worried about particular violations of monotonicity, but there will never be consensus: some audience member is always unmoved by the data on monotonicity.
@@ 2. Setup
Consider a finite population of $n$ individuals subjected to a completely randomized experiment with $n_{1}$ treated units, where $1<n_{1}<n$. That is:
@@
$\theta\in\Theta\equiv\left\{{(n_{at},n_{nt},n_{d},n_{c})\in(\mathbb{N}\cup\left\{{0}\right\})^{4}:n_{at}+n_{nt}+n_{d}+n_{c}=n}\right\}.$
@@
Sometimes, we think of the finite population as being drawn from an infinite superpopulation. To that end, let $\mathcal{P}=\left\{{(p_{at},p_{nt},p_{d},p_{c})\in[0,1]^{4}:p_{at}+p_{nt}+p_{d}+p_{c}=1}\right\}$ be the simplex. For an element $p\in\mathcal{P}$, let the corresponding distribution of $(Y_{T},Y_{U})$ when unit types are drawn i.i.d. from $\mathrm{Multinomial}(1,p)$ be
@@
Whether we prefer to test the superpopulation null that $p\in\mathcal{P}_{0}$ or the design-based null that $\theta_{0}\in\Theta_{0}$ will depend on the application. If the $n$ units are patients in a drug trial drawn randomly from a much larger population of patients with the same condition, then we are likely more interested in whether there are heterogeneous responses to the drug in the super-population of patients, and thus $p\in\mathcal{P}_{0}$ is more relevant than $\theta\in\Theta_{0}$. On the other hand, if the $n$ units are the 50 states, it may be unnatural to imagine the states as sampled from an infinite superpopulation, rendering conceptual issues for the null $p\in\mathcal{P}_{0}$. By contrast, testing $\theta\in\Theta_{0}$ answers the natural question as to whether any of the 50 states have opposite-signed treatment effects. See copas_randomization_1973, reichardt_justifying_1999, and rambachan_design-based_2025, among others, for additional discussion of the relevance of design-based vs. superpopulation based estimands.
@@ 3. Identification
The typical textbook definition of identification states that a parameter is identified if two distinct values of the parameter induce different distributions of the observed data. [^5] For example, the Wikipedia page on identifiability states that a statistical model $P_{\theta}$ is “identifiable if the mapping $\theta\mapsto P_{\theta}$ is one-to-one.” lehmann_theory_1998 equivalently define $\theta$ to be unidentifiable if there exist $\theta_{1}\neq\theta_{2}$ such that $P_{\theta_{1}}=P_{\theta_{2}}$. It is well-known that, from the superpopulation perspective, the population shares $p$ are not point-identified (heckman_making_1997). In particular, any type proportion that violates monotonicity induces data that can be rationalized by some other type proportion that obeys monotonicity. By contrast, as summarized in the following result, the type counts $\theta$ are in fact identified, and thus monotonicity violations are likewise distinguishable from $f_{\theta}$.
@@ 4. Frequentist testing
(Frequentist tests exist) Suppose $n\geq 4$ and $n_{0},n_{1}\geq 2$. Then there exist tests for monotonicity that control size and have power against some alternatives. That is, for any $\alpha\in(0,1)$ there exists a test $\delta:\operatorname{supp}(Y)\to[0,1]$ such that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta}[\delta(Y)]\leq\alpha$ and $E_{\theta_{1}}[\delta(Y)]>\alpha$ for some $\theta_{1}\in\Theta_{1}$. [^6] The proof shows that there is a non-randomized test satisfying the conditions of the proposition for $\alpha$ sufficiently large.
@@
Observe that in Example˜1, the support of $Y$ was $\{(0,1),(1,0)\}$ when there was 1 always-taker and 1 never-taker, but was $\{(0,0),(1,1)\}$ when there was 1 complier and 1 defier. This illustrates that the support of $Y$ may be different when monotonicity holds versus when it is violated. This general idea is the basis of the construction of tests for monotonicity in the proof to Proposition˜4.1. We show that when $n\geq 4$, there always exists an alternative $\theta_{1}$ such that the support of $Y$ under $\theta_{1}$ does not contain the support of $Y$ under $\theta_{0}$ for any $\theta_{0}$ satisfying the null: i.e. $S^{Y}_{\theta_{1}}\not\supseteq S^{Y}_{\theta_{0}}$ for any $\theta_{0}\in\Theta_{0}$, where $S^{Y}_{\theta}$ is the support of $Y$ under $\theta$. [^7] Note that this is not the case in Example 1, because the support of $Y$ when there are two always-takers is $(1,1)\subseteq S^{Y}_{\theta_{1}}=\{(0,0),(1,1)\}$. There are thus no non-trivial tests of monotonicity with $n_{1}=n_{0}=1$. It follows that a test that rejects if and only if $Y\in S^{Y}_{\theta}$ has power of 1 to reject $\theta_{1}$, but only has size $\alpha_{0}=\sup_{\theta\in\Theta_{0}}E_{\theta}[1\{Y\in S^{Y}_{\theta_{1}}\}]<1$. We can then construct a non-trivial randomized test with arbitrary size $\alpha\leq\alpha_{0}$ by rejecting with probability $\alpha/\alpha_{0}>\alpha$ when $Y\in S^{Y}_{\theta_{1}}$.
Although non-trivial tests against monotonicity exist from the design-based perspective, we expect their power to be poor in practice. Our next result shows that any test of monotonicity has trivial power against some alternatives. This implies that there are no consistent tests of monotonicity (that is, tests for which power converges to 1 for all alternatives along a sequence of finite-populations with $n\to\infty$).
@@
(Trivial power for some alternative) Every test of monotonicity has trivial power against some alternatives. That is, suppose that $\delta:\operatorname{supp}(Y)\to[0,1]$ is such that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha$ for some $\alpha\in(0,1)$. Then there exists some $\theta_{1}\in\Theta_{1}$ such that $E_{\theta_{1}}[\delta(Y)]\leq\alpha$.
@@
(Low WAP) Assume $n\geq 4$. Fix any $\theta\in\Theta_{1}$ and let $v=\frac{1}{n}\min(\theta_{c},\theta_{d})\geq 1/n$ be the size of the monotonicity violation. Let $\vartheta\sim\mathrm{Multinomial}(n,\theta/n)$. Consider the weight function $w_{n}(\cdot;\theta)$ over $\Theta_{1}$ defined by the probability mass function of $\vartheta\mid(\vartheta\in\Theta_{1})$. Fix a test $\delta(\cdot)$ such that $\sup_{\theta\in\Theta_{0}}E_{\theta}[\delta(Y)]\leq\alpha$. Then, the weighted average power around $\theta$ is bounded,
$\mathrm{WAP}(\delta;\theta)\equiv\sum_{\vartheta\in\Theta_{1}}E_{\vartheta}[\delta(Y)]w_{n}(\vartheta;\theta)\leq\alpha\frac{1}{1-2(1-v)^{n}+(1-2v)^{n}}\leq 2.51\alpha.$
@@
To illustrate these results, Figure˜1 computes the most powerful 5%-level test for every possible alternative $\theta\in\Theta_{1}$, obtained via linear programming, for $n=30$ and $n_{1}=15$. Perhaps surprisingly, all 4495 alternatives are testable, in the sense that for each alternative, there exists a test targeted to that alternative with power more than 0.05. However, the power for these tests for the targeted alternative tends to be modest: among these tests engineered to maximize power at a given alternative, only 53 of 4495 alternatives have tests with power above 0.15, and all of them have power below 0.31. These tests also tend to have very poor power for nearby alternatives, as suggested by Proposition˜4.3: Only 18 have weighted average power above the nominal threshold $0.05$, and the maximum WAP is a measly $0.0567$. As a specific example, the optimal test against 18 defiers and 12 compliers achieves the maximal power of $0.31$, but this test rejects with probability zero when there are 17 defiers and 13 compliers.
@@
We can obtain even sharper limits on power if we restrict attention to unbiased tests. Recall that a test is unbiased if its power against all alternatives is weakly greater than its size. Proposition˜4.4 below shows that unbiased tests for monotonicity do exist (at least when $n_{1}=n_{0}$). However, Proposition˜4.5, which is a consequence of Proposition˜4.3 above, implies that as the sample size grows large, the power of any unbiased test against any alternative becomes trivial.
@@
(Unbiased tests exist) Suppose $n_{1}=n_{0}\geq 2$. Then there exists a non-trivial unbiased test of monotonicity: for any $\alpha\in(0,1)$, there exists $\delta:\operatorname{supp}(Y)\to[0,1]$ such that $\sup_{\theta\in\Theta_{0}}E_{\theta}[\delta(Y)]\leq\alpha\leq\inf_{\theta\in\Theta_{1}}E_{\theta}[\delta(Y)]$, with $E_{\theta_{1}}[\delta(Y)]>\alpha$ for at least one $\theta_{1}\in\Theta_{1}$.
@@
(Unbiased tests have asymptotically trivial power) Fix $\epsilon>0$ and $n\geq 4$. Fix any $\theta\in\Theta_{1}$ such that $v=\min(\theta_{c},\theta_{d})/n\geq\epsilon$. Let $\delta$ be any unbiased level-$\alpha$ test, i.e. a test satisfying $\sup_{\theta\in\Theta_{0}}E_{\theta}[\delta(Y)]\leq\alpha\leq\inf_{\theta\in\Theta_{1}}E_{\theta}[\delta(Y)].$ Then we have that

### v2

@@ Abstract.
We consider the extent to which we can learn from a completely randomized experiment whether all individuals have treatment effects that are weakly of the same sign, a condition we call monotonicity. From a classical sampling perspective, it is well-known that monotonicity is not falsifiable. By contrast, we show from the design-based perspective—in which the units in the population are fixed and only treatment assignment is stochastic—that the distribution of treatment effects in the finite population (and hence whether monotonicity holds) is formally identified. We argue, however, that the usual definition of identification is unnatural in the design-based setting because it imagines knowing the distribution of outcomes over different treatment assignments for the same units. We thus evaluate the informativeness of the data by the extent to which it enables frequentist testing and Bayesian updating. We show that frequentist tests can have nontrivial power against some alternatives, but power is generically limited. Likewise, we show that there exist (non-degenerate) Bayesian priors that never update about whether monotonicity holds. We conclude that, despite the formal identification result, the ability to learn about monotonicity from data in practice is severely limited.
@@ 1. Introduction
By contrast, the design-based perspective views the $n$ units and their potential outcomes as fixed or conditioned upon, with statistical uncertainty arising only from the stochastic assignment of treatment. The finite population is characterized by $\theta=(\theta_{at},\theta_{nt},\theta_{d},\theta_{c})$, where $\theta_{t}$ is the count of how many of the $n$ units are of type $t$. The design-based monotonicity assumption is then that $\min\{\theta_{c},\theta_{d}\}=0$. [^2] This monotonicity assumption is distinct from the one-sided hypothesis that there are no defiers (or equivalently, no negative treatment effects), which has the (generally testable) implication that the average treatment effect is non-negative. caughey2023randomisation employ randomization inference to construct valid test of the null of no negative treatment effects in a design-based model that also captures non-binary outcomes. In sum, the classical monotonicity assumption asks whether there are any units with opposite-signed treatment effects in the superpopulation from which the sample is drawn, while the design-based version asks whether there are any opposite-signed treatment effects among the $n$ units at hand.
It is well-known that the observable data are always compatible with the null that monotonicity holds in the superpopulation. Suppose that we have a completely randomized experiment in which $n_{1}$ of the $n$ units are randomly assigned to treatment. The researcher then observes $(Y_{T},Y_{U})=(\sum_{i}D_{i}Y_{i},\sum_{i}(1-D_{i})Y_{i})$, summarizing how many units in each treatment group have $Y_{i}=1$. From the superpopulation perspective, we have that $(Y_{T},Y_{U})\sim g_{p}$ (where $g_{p}$ is is the push-forward of $g_{p}^{*}$ under completely random assignment with probability $n_{1}/n$). However, it is straightforward to show that $p$ is not point-identified (see, e.g., heckman_making_1997). In particular, for any $p_{1}$ that violates monotonicity, there exists a $p_{0}$ satisfying monotonicity such that $g_{p_{1}}=g_{p_{0}}$. Hence, from the superpopulation perspective, the monotonicity assumption is not falsifiable. [^3] We note that monotonicity may be testable when combined with other assumptions. For example, there are tests of the joint assumptions of instrument monotonicity, exclusion, and independence (e.g. kitagawa_test_2015). In addition, monotonicity has testable implications for the marginal distributions of treated and control observations when the outcome has more than two levels (angrist1995two; see also Appendix 7.2 of caughey2023randomisation for an application in a design-based framework).
By contrast, we show that from the design-based perspective, the type counts $\theta$ are in fact identified according to the usual technical definition of identification. Let $f_{\theta}$ denote the distribution of the observable data $(Y_{T},Y_{U})$ over repeated assignments of treatment for the fixed finite population. The mapping $\theta\mapsto f_{\theta}$ is one-to-one, so $\theta$ is formally identified. At first blush, the typical definition of identification appears to suggest that the data are informative about whether the design-based version of monotonicity holds.
@@
From the frequentist perspective, we show that there exist tests of the design-based monotonicity assumption with non-trivial power against some alternatives, but we generally expect the power of these tests to be poor. In particular, we show that any test of monotonicity will have trivial or near-trivial power against some alternative: when $n$ is even, any level-$\alpha$ test has power no larger than $\alpha$ against some alternative; when $n$ is odd, worst-case power is bounded above by $\alpha(1+O(2^{-n}))$. Moreover, we show that any test that has power against some alternative $\theta_{1}$ will have poor power at alternatives “close” to $\theta_{1}$. We formalize this by deriving upper bounds on the weighted average power (WAP) of tests in a neighborhood of $\theta_{1}$, where the weights are proportional to the frequencies expected under sampling types from a superpopulation (and thus concentrate around $\theta_{1}$, in terms of relative frequencies, when $n$ is reasonably large). We show that regardless of sample size, the WAP of a size-$\alpha$ test is no greater than $2.51\alpha$; thus, for a 5% test, the WAP is never more than 12.6%. The bound can be even tighter given a particular $n$ and a lower bound on the number of defiers/compliers. For example, if $n=100$ and compliers and defiers are each at least 5% of the finite population, then WAP is never greater than 5.06%. We also derive upper bounds on power for unbiased tests (i.e. those with power at least $\alpha$ for all alternatives). This upper bound converges to $\alpha$ as $n$ grows large, indicating that with large $n$ all unbiased tests have nearly trivial power.
@@
Taken together, our results suggest that frequentist tests provide limited learning about monotonicity. In particular, these tests can only be useful if we are interested in testing one particular alternative and not nearby ones. For example, with $n=30$ and $n_{1}=15$, the highest possible power of a 5% test against any alternative is 31%, which is achieved by a test targeting the alternative with 18 defiers and 12 compliers. However, this test never rejects when there are 17 defiers and 13 compliers. Our theoretical results show the low power of this test against neighboring alternatives is a generic feature of tests against monotonicity. We suspect that in most practical settings, researchers will not be interested in ultra-specific alternatives, and thus such frequentist tests will have relatively little use in practice.
@@
From the Bayesian perspective, we show that there do exist Bayesians who update about the probability that the design-based monotonicity assumption holds, but there also exist Bayesians who never update. The existence of a Bayesian who updates follows immediately from the fact that $\theta$ is identified. Consider a Bayesian with a two-point prior on $\theta_{0}$ satisfying monotonicity and $\theta_{1}$ violating monotonicity. Since $f_{\theta_{0}}\neq f_{\theta_{1}}$, the Bayesian updates about the probability that the null is true. However, we show that there exists a non-trivial prior $\pi$ which never updates on the probability that the null hypothesis holds, i.e. $\pi(\theta\in\Theta_{0}\mid(Y_{U},Y_{T}))=\pi(\theta\in\Theta_{0})\in(0,1)$ ($\pi$-a.s.), where $\Theta_{0}$ denotes the set of parameter values satisfying monotonicity. [^4] Specifically, we show such priors exist when $n$ is even. When $n$ is odd, we show that there exist priors that minimally update, in the sense that the expected absolute difference between the prior and posterior probabilities for $\theta\in\Theta_{0}$ is $O(2^{-n})$; see Proposition A.1. Our results thus suggest that the data may be informative to some audience members, who are worried about particular violations of monotonicity, but there will not be consensus: some audience member is always unmoved by the data on monotonicity. The existence of a Bayesian who doesn’t update also implies that any classifier of whether $\theta\in\Theta_{0}$ does no better than random guessing at some parameter value.
@@ 2. Setup
Consider a finite population of $n$ individuals subjected to a completely randomized experiment with $n_{1}$ treated units, where $0<n_{1}<n$. That is:
@@
$\theta\in\Theta:=\left\{{(n_{at},n_{nt},n_{d},n_{c})\in(\mathbb{N}\cup\left\{{0}\right\})^{4}:n_{at}+n_{nt}+n_{d}+n_{c}=n}\right\}.$
@@
Sometimes, we think of the finite population as being drawn from an infinite superpopulation. To that end, let $\mathcal{P}=\left\{{(p_{at},p_{nt},p_{d},p_{c})\in[0,1]^{4}:p_{at}+p_{nt}+p_{d}+p_{c}=1}\right\}$ be the simplex. For an element $p\in\mathcal{P}$, let the corresponding distribution of $(Y_{T},Y_{U})$ when unit types are drawn i.i.d. from $\mathrm{Multinomial}(n,p)$ be
@@
Whether we prefer to test the superpopulation null that $p\in\mathcal{P}_{0}$ or the design-based null that $\theta_{0}\in\Theta_{0}$ will depend on the application. If the $n$ units are patients in a drug trial drawn randomly from a much larger population of patients with the same condition, then we are likely more interested in whether there are heterogeneous responses to the drug in the super-population of patients, and thus $p\in\mathcal{P}_{0}$ is more relevant than $\theta\in\Theta_{0}$. On the other hand, if the $n$ units are the 50 states, it may be unnatural to imagine the states as sampled from an infinite superpopulation, rendering conceptual issues for the null $p\in\mathcal{P}_{0}$. By contrast, testing $\theta\in\Theta_{0}$ answers the natural question as to whether any of the 50 states have opposite-signed treatment effects. See copas_randomization_1973, reichardt_justifying_1999, and rambachan_design-based_2025, among others, for additional discussion of the relevance of design-based vs. superpopulation-based estimands in general. For our setting specifically, gelman2025russian argues that cases in which decision loss depends on counterfactual outcomes should be analyzed through a framework in which potential outcomes are stochastic.
@@ 3. Identification
The typical textbook definition of identification states that a parameter is (point-)identified if two distinct values of the parameter induce different distributions of the observed data. [^6] For example, the Wikipedia page on identifiability states that a statistical model $P_{\theta}$ is “identifiable if the mapping $\theta\mapsto P_{\theta}$ is one-to-one.” lehmann_theory_1998 equivalently define $\theta$ to be unidentifiable if there exist $\theta_{1}\neq\theta_{2}$ such that $P_{\theta_{1}}=P_{\theta_{2}}$. It is well-known that, from the superpopulation perspective, the population shares $p$ are not point-identified (heckman_making_1997). In particular, any type proportion that violates monotonicity induces data that can be rationalized by some other type proportion that obeys monotonicity. By contrast, as summarized in the following result, the type counts $\theta$ are in fact identified, and thus monotonicity violations are likewise distinguishable from $f_{\theta}$.
@@ 4. Frequentist testing
(Frequentist tests exist) Suppose $n_{0},n_{1}\geq 2$. Then there exist tests for monotonicity that control size and have power against some alternatives. That is, for any $\alpha\in(0,1)$ there exists a test $\delta:\operatorname{supp}(Y)\to[0,1]$ such that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha$ and $E_{\theta_{1}}[\delta(Y)]>\alpha$ for some $\theta_{1}\in\Theta_{1}$. [^7] The proof shows that there is a non-randomized test satisfying the conditions of the proposition for $\alpha$ sufficiently large.
@@
Observe that in Example˜1, the support of $Y$ was $\{(0,1),(1,0)\}$ when there was 1 always-taker and 1 never-taker, but was $\{(0,0),(1,1)\}$ when there was 1 complier and 1 defier. This illustrates that the support of $Y$ may be different when monotonicity holds versus when it is violated. This general idea is the basis of the construction of tests for monotonicity in the proof to Proposition˜4.1. We show that when $n\geq 4$, there always exists an alternative $\theta_{1}$ such that the support of $Y$ under $\theta_{1}$ does not contain the support of $Y$ under $\theta_{0}$ for any $\theta_{0}$ satisfying the null: i.e. $S^{Y}_{\theta_{1}}\not\supseteq S^{Y}_{\theta_{0}}$ for any $\theta_{0}\in\Theta_{0}$, where $S^{Y}_{\theta}$ is the support of $Y$ under $\theta$. [^8] Note that this is not the case in Example 1, because the support of $Y$ when there are two always-takers is $(1,1)\subseteq S^{Y}_{\theta_{1}}=\{(0,0),(1,1)\}$. There are thus no non-trivial tests of monotonicity with $n_{1}=n_{0}=1$. It follows that a test that rejects if and only if $Y\in S^{Y}_{\theta_{1}}$ has power of 1 to reject $\theta_{1}$, but only has size $\alpha_{0}=\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[1\{Y\in S^{Y}_{\theta_{1}}\}]<1$. We can then construct a non-trivial randomized test with arbitrary size $\alpha\leq\alpha_{0}$ by rejecting with probability $\alpha/\alpha_{0}>\alpha$ when $Y\in S^{Y}_{\theta_{1}}$.
Although non-trivial tests against monotonicity exist from the design-based perspective, we expect their power to be poor in practice. Our next result shows that any test of monotonicity has near-trivial power against some alternatives. Specifically, we show that if $n$ is even, there always exists an alternative for which power is no better than size ($\alpha$). For $n$ odd, we show there exists an alternative such that power is bounded above by $\alpha(1+O(2^{-n}))$. [^9] Surprisingly, there is a test for $n=13,n_{1}=6$ whose minimum power over $\Theta_{1}$ is ever-so-slightly larger than $\alpha$. For $\alpha=0.05$, such a test achieves power at least $0.05+2\times 10^{-7}$ over all of $\Theta_{1}$. This implies that there are no consistent tests of monotonicity (that is, tests for which power converges to 1 for all alternatives along a sequence of finite-populations with $n\to\infty$).
@@
(Near-trivial power for some alternative) Suppose that $\delta:\operatorname{supp}(Y)\to[0,1]$ is a level-$\alpha$ test, i.e. $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha$ for some $\alpha\in(0,1)$. If $n$ is even, then $\delta$ has trivial power against some alternative: there exists $\theta_{1}\in\Theta_{1}$ such that $E_{\theta_{1}}[\delta(Y)]\leq\alpha$. If $n$ is odd, then there exists $\theta_{1}\in\Theta_{1}$ such that $E_{\theta_{1}}[\delta(Y)]\leq\alpha\left(1+\frac{1}{2^{n-1}-1}\right)$.
@@
(Low WAP) Assume $n\geq 4$. Fix any $\theta_{1}\in\Theta_{1}$ and let $v=\frac{1}{n}\min(\theta_{1,c},\theta_{1,d})\geq 1/n$ be the size of the monotonicity violation. Let $\vartheta\sim\mathrm{Multinomial}(n,\theta_{1}/n)$. Consider the weight function $w_{n}(\cdot;\theta_{1})$ over $\Theta_{1}$ defined by the probability mass function of $\vartheta\mid(\vartheta\in\Theta_{1})$. Fix a test $\delta(\cdot)$ such that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha$. Then, the weighted average power around $\theta_{1}$ is bounded,
$\mathrm{WAP}(\delta;\theta_{1}):=\sum_{\tilde{\theta}_{1}\in\Theta_{1}}E_{\tilde{\theta}_{1}}[\delta(Y)]w_{n}(\tilde{\theta}_{1};\theta_{1})\leq\alpha\frac{1}{1-2(1-v)^{n}+(1-2v)^{n}}\leq 2.51\alpha.$
@@
To illustrate these results, Fig.˜1 computes the most powerful 5%-level test for every possible alternative $\theta\in\Theta_{1}$, obtained via linear programming, for $n=30$ and $n_{1}=15$. Perhaps surprisingly, all 4495 alternatives are testable, in the sense that for each alternative, there exists a test targeted to that alternative with power more than 0.05. However, the power for these tests for the targeted alternative tends to be modest: among these tests engineered to maximize power at a given alternative, only 53 of 4495 alternatives have tests with power above 0.15, and all of them have power below 0.31. These tests also tend to have very poor power for nearby alternatives, as suggested by Proposition˜4.3: Only 18 have weighted average power above the nominal threshold $0.05$, and the maximum WAP is a measly $0.0567$. As a specific example, the optimal test against 18 defiers and 12 compliers achieves the maximal power of $0.31$, but this test rejects with probability zero when there are 17 defiers and 13 compliers.
@@
We can obtain even sharper limits on power if we restrict attention to unbiased tests. Recall that a test is unbiased if its power against all alternatives is weakly greater than its size. Proposition˜4.4 below shows that unbiased tests for monotonicity do exist (at least when $n_{1}=n_{0}$). However, Proposition˜4.5 implies that as the sample size grows large, the power of any unbiased test against any alternative becomes trivial.
@@
(Unbiased tests exist) Suppose $n_{1}=n_{0}\geq 2$. Then there exists a non-trivial unbiased test of monotonicity: for any $\alpha\in(0,1)$, there exists $\delta:\operatorname{supp}(Y)\to[0,1]$ such that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha\leq\inf_{\theta_{1}\in\Theta_{1}}E_{\theta_{1}}[\delta(Y)]$, with $E_{\theta_{1}}[\delta(Y)]>\alpha$ for at least one $\theta_{1}\in\Theta_{1}$.
@@
(Unbiased tests have asymptotically trivial power) Fix $\epsilon>0$ and $n\geq 4$. Fix any $\theta\in\Theta_{1}$ such that $v=\min(\theta_{c},\theta_{d})/n\geq\epsilon$. Let $\delta$ be any unbiased level-$\alpha$ test, i.e. a test satisfying $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]\leq\alpha\leq\inf_{\theta_{1}\in\Theta_{1}}E_{\theta_{1}}[\delta(Y)].$ Then we have that